# Bias-Corrected Data Synthesis for Imbalanced Learning

### v1

@@ Abstract
Imbalanced data, where the positive samples represent only a small proportion compared to the negative samples, makes it challenging for classification problems to balance the false positive and false negative rates. A common approach to addressing the challenge involves generating synthetic data for the minority group and then training classification models with both observed and synthetic data. However, since the synthetic data depends on the observed data and fails to replicate the original data distribution accurately, prediction accuracy is reduced when the synthetic data is naively treated as the true data. In this paper, we address the bias introduced by synthetic data and provide consistent estimators for this bias by borrowing information from the majority group. We propose a bias correction procedure to mitigate the adverse effects of synthetic data, enhancing prediction accuracy while avoiding overfitting. This procedure is extended to broader scenarios with imbalanced data, such as imbalanced multi-task learning and causal inference. Theoretical properties, including bounds on bias estimation errors and improvements in prediction accuracy, are provided. Simulation results and data analysis on handwritten digit datasets demonstrate the effectiveness of our method.
Keywords: Bias correction; imbalanced classification; oversampling; prediction accuracy; synthetic data.
@@ 1.1 Background
A common strategy for addressing such challenges in imbalanced classification is data augmentation, which aims to rebalance the samples in different classes by artificially modifying or expanding the training dataset. Resampling-based approaches include undersampling – removing samples from the majority class, and oversampling – expanding the minority class. Undersampling techniques, such as Tomek’s links (Tomek, 1976) and cluster centroid (Lemaître et al., 2017), often suffer from information loss due to discarding potentially informative majority samples. In contrast, oversampling is typically preferred, and various methods have been proposed to enrich the minority class. The reweighting procedure, which assigns higher weights to the minority samples, is equivalent to oversampling by replicating the minority samples. While bootstrap (Efron and Tibshirani, 1994) is a widely used resampling method in statistics, its naive application in oversampling may be sensitive to outliers and may introduce variance inflation. Among oversampling methods, the Synthetic Minority Oversampling TEchnique (SMOTE, Chawla et al. (2002)) has been especially influential. SMOTE generates synthetic samples by interpolating between minority samples and has inspired numerous variants, such as Borderline-SMOTE (Han et al., 2005), ADASYN (He et al., 2008), and safe-level-SMOTE (Bunkhumpornpat et al., 2009), which aim to better capture the geometry of the data and concentrate synthetic sample generation near the decision boundary where classification is challenging. For a comprehensive review of resampling techniques in imbalanced settings, see Mohammed et al. (2020).
@@
Beyond empirical success, recent theoretical studies have examined the statistical properties of synthetic procedures and their impact on classification risk. For example, Elreedy et al. (2024) and Sakho et al. (2024) separately derive the probability distribution of SMOTE-generated synthetic samples, with the latter further proving that the synthetic density function vanishes near the boundary of the minority support. Another widely used augmentation method is Mixup (Zhang et al., 2017), which generates new samples by convex combinations of covariates and their labels. Theoretical results for Mixup include robustness against adversarial attacks and improved generalization by reducing overfitting (Zhang et al., 2020), as well as conditions under which Mixup helps reduce calibration errors (Zhang et al., 2022; Naeini et al., 2015).
Since synthetic samples are typically highly dependent on the original training data, it is crucial to carefully handle such a dependence structure. Tian and Shen (2025) propose a partition-based framework in which one subset of data is used to generate synthetic samples, and the other independent subset is used for training. Nevertheless, a fundamental question remains unresolved: Under what conditions do synthetic procedures improve classification, and how can their potential adverse effects be avoided?
1.2 Our Contribution
@@ 2 Methodology
We begin by introducing the setting for binary classification with imbalanced data. Suppose that the training data consist of $n$ independent and identically distributed (i.i.d.) samples $(\bm{X}_{i},Y_{i})_{i=1}^{n}$, where $\bm{X}_{i}\in\mathbb{R}^{d}$ is a $d$-dimensional covariate vector and $Y_{i}\in\{0,1\}$ represents the class label. Assume that $Y_{i}$’s follow a Bernoulli distribution with $\pi_{1}=\mathbb{P}(Y_{i}=1)$ and $\pi_{0}=\mathbb{P}(Y_{i}=0)$. In the imbalanced setting, we assume $0<\pi_{1}\leq 1/2$, so that the class $Y=1$ is underrepresented. For convenience, we refer to $Y=1$ as the minority class and $Y=0$ as the majority class. Let $n_{1}=\sum_{i=1}^{n}Y_{i}$ and $n_{0}=n-n_{1}$ denote the respective sample sizes, with $n_{1}\ll n_{0}$ with high probability. Without loss of generality, we assume that the samples are ordered such that $Y_{1}=\dots=Y_{n_{1}}=1$ and $Y_{n_{1}+1}=\dots=Y_{n}=0$. We also assume that $\bm{X}\mid(Y=1)\sim\mathcal{P}_{1}$ and $\bm{X}\mid(Y=0)\sim\mathcal{P}_{0}$, where $\mathcal{P}_{1},\mathcal{P}_{0}$ represent the class-conditional distributions.
@@ 2.1 Bias Correction with Synthetic Data
Our goal is to build a model to efficiently predict $Y_{n+1}$ based on the new covariate vector $\bm{X}_{n+1}$. For probability prediction function $f\in\mathcal{F}$, where $\mathcal{F}=\{f:\mathbb{R}^{d}\to[0,1]\}$, denote the corresponding loss function as a binary cross-entropy loss, for example, $\ell_{f}(\bm{X},Y)=-Y\log f(\bm{X})-(1-Y)\log(1-f(\bm{X}))$. With the raw data $(\bm{X}_{i},Y_{i})_{i=1}^{n}$, we can just train the prediction function by minimizing the empirical loss function $L^{\rm raw}$:
@@
In the case that $n_{1}\ll n_{0}$, a trivial guess that all samples are from the majority group will result in accuracy as high as $n_{0}/n$, which is close to $1$, but it does not provide information from the data. To deal with this problem, an intuitive way is to make the data balanced by adding synthetic data samples to the minority group. Assume that we have $\tilde{n}_{1}$ synthetic samples for the minority group: $(\tilde{\bm{X}}^{(1)}_{i},\tilde{Y}_{i}^{(1)})_{i=1}^{\tilde{n}_{1}}$, where $\tilde{Y}_{i}^{(1)}=1$ for $i=1,\ldots,\tilde{n}_{1}$. By equally treating the synthetic and raw samples, we can run the algorithm by minimizing the synthetic-augmented loss function $L^{\rm syn}$:
@@
Figure 1: A pictorial illustration of imbalanced learning based on raw data, synthetic augmentation and bias correction.
Consider $n_{1}^{*}$ which satisfies $n_{1}+n_{1}^{*}\approx n_{0}$. Imagine that we have $n_{1}^{*}$ unobserved samples from the minority group $(\bm{X}_{i}^{*},Y_{i}^{*})_{i=1}^{n_{1}^{*}}$ where $Y_{i}^{*}=1$ and $\bm{X}_{i}^{*}\mid(Y_{i}^{*}=1)\sim\mathcal{P}_{1}$. By introducing the unobserved minority samples, the total dataset $(\bm{X}_{i},Y_{i})_{i=1}^{n}$ and $(\bm{X}_{i}^{*},Y_{i}^{*})_{i=1}^{n_{1}^{*}}$ is roughly balanced. Denote the sample bias caused by the synthetic data as
$\hat{\Delta}_{1}=\frac{1}{n_{1}^{*}}\sum_{i=1}^{n_{1}^{*}}\ell_{f}(\bm{X}_{i}^{*},Y_{i}^{*})-\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},\tilde{Y}_{i}^{(1)}).$ (2.4)
@@
$=$ $\frac{1}{n+n_{1}^{*}}\left\{\sum_{i=1}^{n}\ell_{f}(\bm{X}_{i},Y_{i})+n_{1}^{*}\cdot\left(\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},\tilde{Y}_{i}^{(1)})+\hat{\Delta}_{1}\right)\right\}.$ (2.5)
Since $(\bm{X}_{i}^{*},Y_{i}^{*})_{i=1}^{n_{1}^{*}}$ are not observable, it is impossible to calculate $\hat{\Delta}_{1}$ with the observed data. We can estimate the bias from the available data in the majority group as follows. First, randomly partition the majority indices into generation subgroup $\mathcal{S}_{0g}$ and correction subgroup $\mathcal{S}_{0c}$ with corresponding sizes $n_{0g}$ and $n_{0c}$, respectively. Next, using samples from the generation subgroup $(\bm{X}_{i},Y_{i})_{i\in\mathcal{S}_{0g}}$, generate $\tilde{n}_{0}$ synthetic samples $(\tilde{\bm{X}}_{i}^{(0)},\tilde{Y}_{i}^{(0)})_{i=1}^{\tilde{n}_{0}}$ by the same synthetic generator, where $\tilde{Y}_{i}^{(0)}=0$ for $i=1,\ldots,\tilde{n}_{0}$. Then consider the population bias of the loss function from the true majority distribution $\mathcal{P}_{0}$ to the synthetic majority distribution $\tilde{\mathcal{P}}_{0}$ by
@@
$\hat{\Delta}_{0}=\frac{1}{n_{0c}}\sum_{i\in\mathcal{S}_{0c}}\ell_{f}(\bm{X}_{i},Y_{i})-\frac{1}{\tilde{n}_{0}}\sum_{i=1}^{\tilde{n}_{0}}\ell_{f}(\tilde{\bm{X}}_{i}^{(0)},\tilde{Y}_{i}^{(0)}).$ (2.7)
Note that all elements for calculating $\hat{\Delta}_{0}$ are available and $\hat{\Delta}_{0}$ is constructed the same way as $\hat{\Delta}_{1}$. Suppose that the transformation from $\mathcal{P}_{0}$ to $\mathcal{P}_{1}$ is captured by a measurable function $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$, and assume that the transformation property is well maintained by the corresponding synthetic distributions. This assumption is formally elaborated in Section 3.1. Given the above properties, it is possible to estimate the unobservable $\hat{\Delta}_{1}$ using the data-driven estimator $\hat{\Delta}_{0}$. Thus, by modifying the balanced loss $L^{\rm bal}(f)$ in (2.5), we propose the following bias-correction loss function $L^{\rm bc}(f)$:
$\begin{split}L^{\rm bc}(f)=\frac{1}{n+\tilde{n}_{1}}\bigg[\sum_{i=1}^{n}\ell_{f}(\bm{X}_{i},Y_{i})+\tilde{n}_{1}\cdot\bigg\{\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},\tilde{Y}_{i}^{(1)})+\hat{\Delta}_{0}\bigg\}\bigg].\end{split}$ (2.8)
The majority bias correction term $\hat{\Delta}_{0}$ captures the loss function bias induced by the discrepancy between the true and synthetic distributions from the majority group. Under mild assumptions, $\hat{\Delta}_{0}$ is a good representation of $\hat{\Delta}_{1}$, the loss function bias from the minority group, up to a fixed bias-transfer error and sampling fluctuations, as illustrated in Section 3.1. With this property, the term $\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},\tilde{Y}_{i}^{(1)})+\hat{\Delta}_{0}$ is regarded as the average loss from the unobserved minority samples after correcting the synthetic bias. Consequently, the bias-corrected loss $L^{\rm bc}$ represents a valid average loss from a roughly balanced dataset. Finally, we can find the prediction function by minimizing the bias-corrected loss:
$\hat{f}^{\rm bc}=\underset{f\in\mathcal{F}}{\arg\min}\>L^{\rm bc}(f).$ (2.9)
@@
1: Minority augmentation: Generate $\tilde{n}_{1}$ synthetic minority samples
$(\tilde{\bm{X}}_{i}^{(1)})_{i=1}^{\tilde{n}_{1}}\leftarrow\mathcal{G}\big((\bm{X}_{i})_{Y_{i}=1}\big).$
@@
3: Generate $\tilde{n}_{0}$ synthetic majority samples
$(\tilde{\bm{X}}_{i}^{(0)})_{i=1}^{\tilde{n}_{0}}\leftarrow\mathcal{G}\big((\bm{X}_{i})_{i\in\mathcal{S}_{0g}}\big).$
4: Compute the empirical majority bias $\hat{\Delta}_{0}$ according to Equation (2.7).
@@
6: Obtain the predictor by
$\hat{f}^{\rm bc}=\underset{f\in\mathcal{F}}{\arg\min}\>L^{\rm bc}(f).$
6: The prediction function $\hat{f}^{\rm bc}:\mathbb{R}^{d}\to(0,1)$.

### v2

@@ Abstract
Imbalanced data, where the positive samples represent only a small proportion compared to the negative samples, makes it challenging for classification problems to balance the false positive and false negative rates. A common approach to addressing the challenge involves generating synthetic data for the minority group and then training classification models with both observed and synthetic data. However, since the synthetic data depends on the observed data and fails to replicate the original data distribution accurately, prediction accuracy is reduced when the synthetic data is naïvely treated as the true data. In this paper, we address the bias introduced by synthetic data and provide consistent estimators for this bias by borrowing information from the majority group. We propose a bias correction procedure to mitigate the adverse effects of synthetic data, enhancing prediction accuracy while avoiding overfitting. This procedure is extended to broader scenarios with imbalanced data, such as imbalanced multi-task learning and causal inference. Theoretical properties, including bounds on bias estimation errors and improvements in prediction accuracy, are provided. Simulation results and data analysis on handwritten digit datasets demonstrate the effectiveness of our method.
Keywords: Bias correction; Empirical risk minimization; Imbalanced classification; Oversampling; Synthetic data.
@@ 1.1 Background
A common strategy for addressing such challenges in imbalanced classification is data augmentation, which aims to rebalance the samples in different classes by artificially modifying or expanding the training dataset. Resampling-based approaches include undersampling (removing samples from the majority class) and oversampling (expanding the minority class). Undersampling techniques, such as Tomek’s links (Tomek, 1976) and cluster centroid (Lemaître et al., 2017), often suffer from information loss due to discarding potentially informative majority samples. In contrast, oversampling is typically preferred, and various methods have been proposed to enrich the minority class. The reweighting procedure, which assigns higher weights to the minority samples, is equivalent to oversampling by replicating the minority samples. While bootstrap (Efron and Tibshirani, 1994) is a widely used resampling method in statistics, its naïve application in oversampling may be sensitive to outliers and may introduce variance inflation. The Gaussian mixture model (McLachlan and Peel, 2000) is a classical tool for modeling heterogeneous populations through parametric mixture distributions. Perturbed sampling techniques (He and Garcia, 2009) have been widely used to address class imbalance by modifying the sampling mechanism across classes. Both approaches provide convenient modeling abstractions, though they rely on simplified assumptions and may not fully capture the complexity of real-world data.
Among oversampling methods, the Synthetic Minority Oversampling TEchnique (SMOTE, Chawla et al. (2002)) has been especially influential. SMOTE generates synthetic samples by interpolating between minority samples and has inspired numerous variants, such as Borderline-SMOTE (Han et al., 2005), ADASYN (He et al., 2008), and safe-level-SMOTE (Bunkhumpornpat et al., 2009), which aim to better capture the geometry of the data and concentrate synthetic sample generation near the decision boundary where classification is challenging. For a comprehensive review of resampling techniques in imbalanced settings, see Mohammed et al. (2020).
@@
Beyond empirical success, recent theoretical studies have examined the statistical properties of synthetic procedures and their impact on classification risk. For example, Elreedy et al. (2024) and Sakho et al. (2024) separately derive the probability distribution of SMOTE-generated synthetic samples, with the latter further proving that the synthetic density function vanishes near the boundary of the minority support. Another widely used augmentation method is Mixup (Zhang et al., 2017), which generates new samples by convex combinations of covariates and their labels. In Xia and Klusowski (2026), imbalanced classification is viewed through a transfer-learning lens under label shift, and regimes are identified in which oversampling can be detrimental. From a learning-theoretic perspective, uniform concentration bounds are established in Ahmad et al. (2025) that link the empirical risk computed on synthetic samples to the true minority risk, yielding excess-risk guarantees for kernel-based classifiers. Theoretical results for Mixup include robustness against adversarial attacks and improved generalization by reducing overfitting (Zhang et al., 2020), as well as conditions under which Mixup helps reduce calibration errors (Naeini et al., 2015; Zhang et al., 2022).
Some recent work suggests that resampling/upsampling does not always improve performance under class imbalance and can even be counterproductive in certain prediction settings. Large empirical comparisons find that the benefits of common sampling strategies are highly dataset- and model-dependent, with no uniformly best choice across imbalance regimes (Newaz et al., 2022). In clinical risk prediction with logistic regression, it is shown that oversampling and SMOTE can substantially degrade probability calibration without improving discrimination, and that comparable operating points can often be achieved via threshold adjustment (Van den Goorbergh et al., 2022). Assunção et al. (2024) argue, under a calibration-centric perspective, that adjusting the decision cutoff may recover much of the benefit typically attributed to augmentation. Motivated by the dependence introduced when synthetic samples are generated from the training data, Tian and Shen (2025) propose a sample-splitting framework in which one subset is used to generate synthetic samples and an independent subset is used for training. Nevertheless, a fundamental question remains: under what conditions do synthetic procedures improve classification, and how can potential adverse effects be avoided? In this work, we aim to address this question by developing a principled framework to analyze and correct the distributional discrepancy introduced by synthetic data.
1.2 Our Contributions
@@ 2 Methodology
We begin by introducing the setting for binary classification with imbalanced data. Suppose that the training data consist of $n$ independent and identically distributed (i.i.d.) samples $(\bm{X}_{i},Y_{i})_{i=1}^{n}$, where $\bm{X}_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ is a $d$-dimensional covariate vector and $Y_{i}\in\{0,1\}$ represents the class label. Assume that $Y_{i}$’s follow a Bernoulli distribution with $\pi_{1}=\mathbb{P}(Y_{i}=1)$ and $\pi_{0}=\mathbb{P}(Y_{i}=0)$. In the imbalanced setting, we assume $0<\pi_{1}<1/2$, so that the class $Y=1$ is underrepresented. For convenience, we refer to $Y=1$ as the minority class and $Y=0$ as the majority class. Let $n_{1}=\sum_{i=1}^{n}Y_{i}$ and $n_{0}=n-n_{1}$ denote the respective sample sizes, with $n_{1}\ll n_{0}$ with high probability. Without loss of generality, we assume that the samples are ordered such that $Y_{1}=\dots=Y_{n_{1}}=1$ and $Y_{n_{1}+1}=\dots=Y_{n}=0$. We also assume that $\bm{X}\mid(Y=1)\sim\mathcal{P}_{1}$ and $\bm{X}\mid(Y=0)\sim\mathcal{P}_{0}$, where $\mathcal{P}_{1},\mathcal{P}_{0}$ represent the class-conditional distributions.
@@ 2.1 Bias Correction with Synthetic Data
Our goal is to learn a prediction function $f:\mathcal{X}\to(0,1)$ that provides accurate probability estimates for $Y_{n+1}$ given a new covariate vector $\bm{X}_{n+1}$. Our framework focuses on risk minimization with respect to a general loss function and does not impose a specific classification or decision rule. In particular, the learned function $f$ can be combined with any downstream classification rule depending on the application, such as thresholding or cost-sensitive decisions. One widely used example is to obtain the prediction via $\widehat{Y}_{n+1}=I(f(\bm{X}_{n+1})\geq 1/2)$ where $I(\cdot)$ denotes the indicator function taking values in $\{0,1\}$. Considering the family of prediction functions $\mathcal{F}=\{f:\mathbb{R}^{d}\to(0,1)\}$, denote the corresponding loss function as a binary cross-entropy loss, for example, $\ell_{f}(\bm{X},Y)=-Y\log f(\bm{X})-(1-Y)\log(1-f(\bm{X}))$. Throughout this paper, the proposed correction is formulated at the loss and risk level and is agnostic to the specific classification architecture. Our framework applies to classifiers $f\in\mathcal{F}$ trained via empirical risk minimization, including (but not limited to) logistic regression and neural networks with cross-entropy loss, as well as margin-based methods such as calibrated support vector machines and boosting.
With the raw data $(\bm{X}_{i},Y_{i})_{i=1}^{n}$, we can simply train the prediction function by minimizing the empirical loss function $L^{\rm raw}$:
@@
In the case that $n_{1}\ll n_{0}$, a trivial guess that all samples are from the majority group will result in accuracy as high as $n_{0}/n$, which is close to $1$, but it does not provide information from the data. To deal with this problem, an intuitive way is to make the data balanced by adding synthetic data samples to the minority group. Assume that we have $\tilde{n}_{1}$ synthetic samples for the minority group: $(\tilde{\bm{X}}^{(1)}_{i},\tilde{Y}_{i}^{(1)})_{i=1}^{\tilde{n}_{1}}$, where $\tilde{Y}_{i}^{(1)}=1$ for $i=1,\ldots,\tilde{n}_{1}$ with $\tilde{n}_{1}\approx n_{0}-n_{1}$. By equally treating the synthetic and raw samples, we can run the algorithm by minimizing the synthetic-augmented loss function $L^{\rm syn}$:
@@
Figure 1: Imbalanced learning based on raw data, synthetic augmentation and bias correction.
Since the sample size in the minority class is limited, and the synthetic samples $(\tilde{\bm{X}}_{i}^{(1)})_{i=1}^{\tilde{n}_{1}}$ are generated conditional on the observed samples $(\bm{X}_{i})_{i=1}^{n_{1}}$, a direct comparison between the empirical losses $n_{1}^{-1}\sum_{i=1}^{n_{1}}\ell_{f}(\bm{X}_{i},1)$ and $\tilde{n}_{1}^{-1}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}},1)$ leads to variance inflation and induces a complicated dependence structure in the resulting difference. To overcome the two challenges, consider $n_{1}^{*}$ satisfying $n_{1}^{*}\approx n_{0}-n_{1}$. Imagine that we have $n_{1}^{*}$ i.i.d. unobserved samples from the minority group $(\bm{X}_{i}^{*},Y_{i}^{*})_{i=1}^{n_{1}^{*}}$ independent of the observed data, where $Y_{i}^{*}=1$ and $\bm{X}_{i}^{*}\mid(Y_{i}^{*}=1)\sim\mathcal{P}_{1}$. By introducing the unobserved minority samples, the total dataset $(\bm{X}_{i},Y_{i})_{i=1}^{n}$ and $(\bm{X}_{i}^{*},Y_{i}^{*})_{i=1}^{n_{1}^{*}}$ is roughly balanced. Denote the sample bias caused by the synthetic data as
$\widehat{\Delta}_{1}=\frac{1}{n_{1}^{*}}\sum_{i=1}^{n_{1}^{*}}\ell_{f}(\bm{X}_{i}^{*},Y_{i}^{*})-\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},\tilde{Y}_{i}^{(1)}).$ (2.4)
@@
$=$ $\frac{1}{n+n_{1}^{*}}\left\{\sum_{i=1}^{n}\ell_{f}(\bm{X}_{i},Y_{i})+n_{1}^{*}\cdot\left(\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},\tilde{Y}_{i}^{(1)})+\widehat{\Delta}_{1}\right)\right\}.$ (2.5)
Since $(\bm{X}_{i}^{*},Y_{i}^{*})_{i=1}^{n_{1}^{*}}$ are not observable, it is impossible to calculate $\widehat{\Delta}_{1}$ with the raw data. As formalized through explicit assumptions introduced later, we can estimate the bias from the available data in the majority group as follows. First, randomly partition the majority indices into a generation subgroup $\mathcal{S}_{0g}$ and a correction subgroup $\mathcal{S}_{0c}$ with corresponding sizes $n_{0g}$ and $n_{0c}$, respectively. In practice, we recommand chossing $n_{0g}\approx n_{1}$ and therefore $n_{0c}\approx n_{0}-n_{1}$. Next, using samples from the generation subgroup $(\bm{X}_{i},Y_{i})_{i\in\mathcal{S}_{0g}}$, generate $\tilde{n}_{0}$ synthetic samples $(\tilde{\bm{X}}_{i}^{(0)},\tilde{Y}_{i}^{(0)})_{i=1}^{\tilde{n}_{0}}$ by the same synthetic generator, where $\tilde{n}_{0}\approx n_{0}-n_{1}$ and $\tilde{Y}_{i}^{(0)}=0$ for $i=1,\ldots,\tilde{n}_{0}$. Then consider the population bias of the loss function from the true majority distribution $\mathcal{P}_{0}$ to the synthetic majority distribution $\tilde{\mathcal{P}}_{0}$ by
@@
$\widehat{\Delta}_{0}=\frac{1}{n_{0c}}\sum_{i\in\mathcal{S}_{0c}}\ell_{f}(\bm{X}_{i},Y_{i})-\frac{1}{\tilde{n}_{0}}\sum_{i=1}^{\tilde{n}_{0}}\ell_{f}(\tilde{\bm{X}}_{i}^{(0)},\tilde{Y}_{i}^{(0)}).$ (2.7)
Note that all elements for calculating $\widehat{\Delta}_{0}$ are available, and $\widehat{\Delta}_{0}$ is constructed the same way as $\widehat{\Delta}_{1}$. Suppose there exists a measurable mapping $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that the minority distribution $\mathcal{P}_{1}$ is well approximated by the pushforward measure $(\mathcal{P}_{0})_{\#T}$, up to a controlled distributional error. Furthermore, assume that this transformation structure is approximately preserved under the corresponding synthetic distributions. This assumption is formally elaborated in Section 3.1. Given the above properties, it is possible to estimate the unobservable $\widehat{\Delta}_{1}$ using the data-driven estimator $\widehat{\Delta}_{0}$. To obtain a computable surrogate of the balanced loss, we further assume that the number of unobserved minority samples $n_{1}^{*}$ is matched by the number of generated synthetic minority samples, i.e., $n_{1}^{*}\approx\tilde{n}_{1}$. Thus, by modifying the balanced loss $L^{\rm bal}(f)$ in (2.5), we propose the following bias-correction loss function $L^{\rm bc}(f)$:
$\begin{split}L^{\rm bc}(f)=\frac{1}{n+\tilde{n}_{1}}\bigg[\sum_{i=1}^{n}\ell_{f}(\bm{X}_{i},Y_{i})+\tilde{n}_{1}\cdot\bigg\{\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},\tilde{Y}_{i}^{(1)})+\widehat{\Delta}_{0}\bigg\}\bigg].\end{split}$ (2.8)
The majority bias correction term $\widehat{\Delta}_{0}$ captures the loss function bias induced by the discrepancy between the true and synthetic distributions from the majority group. Under mild assumptions, $\widehat{\Delta}_{0}$ is a good representation of $\widehat{\Delta}_{1}$, the loss function bias from the minority group, up to a fixed bias-transfer error and sampling fluctuations, as illustrated in Section 3.1. With this property, the term $\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},\tilde{Y}_{i}^{(1)})+\widehat{\Delta}_{0}$ is regarded as the average loss from the unobserved minority samples after correcting the synthetic bias. Consequently, the bias-corrected loss $L^{\rm bc}$ represents a valid average loss from a roughly balanced dataset. Finally, we can find the prediction function by minimizing the bias-corrected loss:
$\widehat{f}^{\rm bc}=\underset{f\in\mathcal{F}}{\arg\min}\>L^{\rm bc}(f).$ (2.9)
@@
1: Minority augmentation: Generate $\tilde{n}_{1}$ synthetic minority samples $(\tilde{\bm{X}}_{i}^{(1)})_{i=1}^{\tilde{n}_{1}}$ via generator $\mathcal{G}$ trained on the minority group $(\bm{X}_{i})_{Y_{i}=1}$.
@@
3: Generate $\tilde{n}_{0}$ synthetic majority samples $(\tilde{\bm{X}}_{i}^{(0)})_{i=1}^{\tilde{n}_{0}}$ via generator $\mathcal{G}$ trained on the generation set $(\bm{X}_{i})_{i\in\mathcal{S}_{0g}}$.
4: Compute the empirical majority bias $\widehat{\Delta}_{0}$ according to Equation (2.7).
@@
6: Obtain the predictor $\widehat{f}^{\rm bc}$ by (2.9).
6: The prediction function $\widehat{f}^{\rm bc}:\mathbb{R}^{d}\to(0,1)$.